# Holistic Evaluation for Interleaved Text-and-Image Generation

### v1

@@ Abstract
Interleaved text-and-image generation has been an intriguing research direction, where the models are required to generate both images and text pieces in an arbitrary order. Despite the emerging advancements in interleaved generation, the progress in its evaluation still significantly lags behind. Existing evaluation benchmarks do not support arbitrarily interleaved images and text for both inputs and outputs, and they only cover a limited number of domains and use cases. Also, current works predominantly use similarity-based metrics which fall short in assessing the quality in open-ended scenarios. To this end, we introduce InterleavedBench, the first benchmark carefully curated for the evaluation of interleaved text-and-image generation. InterleavedBench features a rich array of tasks to cover diverse real-world use cases. In addition, we present InterleavedEval, a strong reference-free metric powered by GPT-4o to deliver accurate and explainable evaluation. We carefully define five essential evaluation aspects for InterleavedEval, including text quality, perceptual quality, image coherence, text-image coherence, and helpfulness, to ensure a comprehensive and fine-grained assessment. Through extensive experiments and rigorous human evaluation, we show that our benchmark and metric can effectively evaluate the existing models with a strong correlation with human judgments surpassing previous reference-based metrics. We also provide substantial findings and insights to foster future research in interleaved generation and its evaluation. [^1] The source code and datasets will be publicly available for research purposes.
@@ 1 Introduction
Multimodal learning has been a rapidly developing research field given the recent advancements in Large Multimodal Models (LMMs) Xu et al. (2023); Dai et al. (2023); Liu et al. (2023). While these models can perform diverse tasks such as detailed image description and visual question answering, the outputs are limited to the text-only format, which hinders their broader applications. More recently, there has been a growing focus on enhancing LMMs with the capability of interleaved generation, i.e., generating multimodal content that seamlessly integrates both text and one or multiple images Koh et al. (2023); Dong et al. (2024); Sun et al. (2023b, a). This opens new avenues for applications in diverse challenging scenarios, such as creative content generation Anantrasirichai and Bull (2022), visual storytelling Huang et al. (2016); Lukin et al. (2018), and multimodal script generation Yang et al. (2021); Qi et al. (2024).
@@ 2 Related Work
Evaluating multimodal generation tasks presents unique challenges due to the inherent complexity of assessing both textual and visual components simultaneously. Traditional metrics for text generation, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005), fall short when applied to multimodal outputs as they fail to capture the visual quality and coherence with textual content. Similarly, visual generation metrics like FID Heusel et al. (2017) and IS Salimans et al. (2016) are inadequate for evaluating the textual elements accompanying the images. To address this, recent studies have employed multimodal metrics such as CLIPScore Hessel et al. (2021), which leverages the alignment capabilities of the CLIP model to measure the similarity between generated images and their corresponding textual descriptions. However, CLIPScore can only measure the alignment between text and images, which is not sufficient to evaluate the quality of generated output comprehensively. Moreover, human evaluations, although more reliable, are resource-intensive and cannot be scalable. In terms of evaluation benchmarks in multimodal learning, existing works mostly focus on evaluating the tasks with single-modality output Jiang et al. (2024), such as, conditional text-to-image generation Chen et al. (2024); Ku et al. (2024), where the primary focus is solely the quality of generated images. Our InterleavedBench benchmark introduces a novel approach to evaluate interleaved text-and-image generation by incorporating multiple aspects of quality assessment, thus providing a more nuanced and holistic evaluation framework.
@@ 3.1 Dataset Curation Process
MagicBrush Zhang et al. (2023) No Single No Single
@@
Collection of Context-based Subset
Firstly, we collect the source data of the context-based subset from existing academic datasets or web resources. Specifically, we collect the data of multimodal script generation from WikiHow Yang et al. (2021), visual story completion from VIST Huang et al. (2016), activity generation from the dense captions and the extracted video frames in ActivityNet Captions Krishna et al. (2017), sequential image editing from MagicBrush Zhang et al. (2023), and multi-concept image composition from CustomDiffusion Kumari et al. (2023a). For web resources, we apply an automatic data filtering pipeline to discard the samples with poor quality to obtain a small set of source data. We detail our data filtering pipeline in Appendix A. Secondly, after collecting the source data (either from academic benchmarks or web resources), we then apply a human selection process to manually select the samples based on data quality and diversity (i.e., avoiding selecting similar samples). Finally, we ask human experts to annotate an instruction $I$ for each sample based on the collected content. We include the details of the data selection and instruction annotation process in Appendix A. For the samples that are originally interleaved articles, we pick the first $k$ images and their associated text as the context $\mathcal{C}$ for the input. $k$ is randomly sampled for each example and ranges from 1 to the maximum number of images minus 1 since we need to ensure the output contains at least one image. The rest of the images and text are used as the gold reference.

### v2

@@ Abstract
Interleaved text-and-image generation has been an intriguing research direction, where the models are required to generate both images and text pieces in an arbitrary order. Despite the emerging advancements in interleaved generation, the progress in its evaluation still significantly lags behind. Existing evaluation benchmarks do not support arbitrarily interleaved images and text for both inputs and outputs, and they only cover a limited number of domains and use cases. Also, current works predominantly use similarity-based metrics which fall short in assessing the quality in open-ended scenarios. To this end, we introduce InterleavedBench, the first benchmark carefully curated for the evaluation of interleaved text-and-image generation. InterleavedBench features a rich array of tasks to cover diverse real-world use cases. In addition, we present InterleavedEval, a strong reference-free metric powered by GPT-4o to deliver accurate and explainable evaluation. We carefully define five essential evaluation aspects for InterleavedEval, including text quality, perceptual quality, image coherence, text-image coherence, and helpfulness, to ensure a comprehensive and fine-grained assessment. Through extensive experiments and rigorous human evaluation, we show that our benchmark and metric can effectively evaluate the existing models with a strong correlation with human judgments surpassing previous reference-based metrics. We also provide substantial findings and insights to foster future research in interleaved generation and its evaluation. [^1] The source code and datasets are publicly available at https://huggingface.co/mqliu/InterleavedBench for research purposes.
@@
https://vt-nlp.github.io/InterleavedEval/
Figure 1: Overview of our InterleavedBench, a comprehensive benchmark that covers 10 diverse use cases for interleaved text-and-image generation, and the evaluation results of InterleavedEval based on GPT-4o.
@@ 1 Introduction
Multimodal learning has been a rapidly developing research field given the recent advancements in Large Multimodal Models (LMMs) Xu et al. (2023); Dai et al. (2023); Liu et al. (2023a). While these models can perform diverse tasks such as detailed image description and visual question answering, the outputs are limited to the text-only format, which hinders their broader applications. More recently, there has been a growing focus on enhancing LMMs with the capability of interleaved generation, i.e., generating multimodal content that seamlessly integrates both text and one or multiple images Koh et al. (2023); Dong et al. (2024); Sun et al. (2023b, a). This opens new avenues for applications in diverse challenging scenarios, such as creative content generation Anantrasirichai and Bull (2022), visual storytelling Huang et al. (2016); Lukin et al. (2018), and multimodal script generation Yang et al. (2021); Qi et al. (2024).
@@ 2 Related Work
Evaluating multimodal generation tasks presents unique challenges due to the inherent complexity of assessing both textual and visual components simultaneously. Existing metrics for text generation, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and LLM-based evaluators Zhong et al. (2022); Liu et al. (2023b, 2024), fall short when applied to multimodal outputs as they fail to capture the visual quality and coherence with textual content. Similarly, visual generation metrics like FID Heusel et al. (2017) and IS Salimans et al. (2016) are inadequate for evaluating the textual elements accompanying the images. To address this, recent studies have employed multimodal metrics Zhang et al. (2023b); Ku et al. (2023), such as CLIPScore Hessel et al. (2021), which leverages the alignment capabilities of the CLIP model to measure the similarity between generated images and their corresponding textual descriptions. However, CLIPScore can only measure the alignment between text and images, which is not sufficient to evaluate the quality of generated output comprehensively. Moreover, human evaluations, although more reliable, are resource-intensive and cannot be scalable. In terms of evaluation benchmarks in multimodal learning, existing works mostly focus on evaluating the tasks with single-modality output Fu et al. (2024); Li et al. (2024); Lu et al. (2024); Wang et al. (2024), such as conditional text-to-image generation Chen et al. (2024); Ku et al. (2024), where the primary focus is solely the quality of generated images. Our InterleavedBench benchmark introduces a novel approach to evaluate interleaved text-and-image generation by incorporating multiple aspects of quality assessment, thus providing a more nuanced and holistic evaluation framework.
@@ 3.1 Dataset Curation Process
Collection of Context-based Subset
Firstly, we collect the source data of the context-based subset from existing academic datasets or web resources. Specifically, we collect the data of multimodal script generation from WikiHow Yang et al. (2021), visual story completion from VIST Huang et al. (2016), activity generation from the dense captions and the extracted video frames in ActivityNet Captions Krishna et al. (2017), sequential image editing from MagicBrush Zhang et al. (2023a), and multi-concept image composition from CustomDiffusion Kumari et al. (2023a). For web resources, we apply an automatic data filtering pipeline to discard the samples with poor quality to obtain a small set of source data. We detail our data filtering pipeline in Appendix A. Secondly, after collecting the source data (either from academic benchmarks or web resources), we then apply a human selection process to manually select the samples based on data quality and diversity (i.e., avoiding selecting similar samples). Finally, we ask human experts to annotate an instruction $I$ for each sample based on the collected content. We include the details of the data selection and instruction annotation process in Appendix A. For the samples that are originally interleaved articles, we pick the first $k$ images and their associated text as the context $\mathcal{C}$ for the input. $k$ is randomly sampled for each example and ranges from 1 to the maximum number of images minus 1 since we need to ensure the output contains at least one image. The rest of the images and text are used as the gold reference.
@@
MagicBrush Zhang et al. (2023a) No Single No Single